OpenStudy (anonymous):

HELP PLEASE! Will medal and fan! Which is equal to 6sqrt32^3 By 6 sqrt I mean the root is 6, not 6 * sqrt 32^2. I'm in logarithms and I checked the lesson.. it's not there D: 4√2 4√3 16√2 16√3

OpenStudy (anonymous):

All those question symbols are actually the sqrt symbol

OpenStudy (lyrae):

\(\large \sqrt[6]{32^3}\) ? Rember that: \[\large \sqrt[n]{a} = a^{\frac{ 1 }{ n }}\]and\[\large (a^n)^k = a^{n \times k}\]

OpenStudy (anonymous):

Ah.. I don't quite know what that means. The thing is this never showed up in my courses. However, my teacher made a recording for the test and I found a question like this in the recording. I got my answer, but thanks for trying to help!

OpenStudy (lyrae):

Ok! Good luck with your studies :)

OpenStudy (anonymous):

Thanks :)